Find the component along x and y axes of a vector 20 unit long when it`s make an angles of 30° with the x axis

Grade:11

2 Answers

Arun
6 years ago
Vector length = 20 unitAngle by x axis = 30°Now component of vector in the directionof x axis = 20 cos 30°= 20* (3^1/2)/ 2 = 20 *1.732/2= 17.32 unitNow component of 20 unit vector in the direction of y axis= 20 sin30= 20 *1/2= 10 unit
Shailendra Kumar Sharma
6 years ago
The vector makes an angle of 30 degree from X axis soX component of vector be 20cos30= 20*(1.732/2) =17.33Whereas y component will be 20sin30 = 10